# Simulation studies on compensation for space-charge-induced half-integer and 3rd-order resonance crossing in HIAF-BRing

### v1

@@ Abstract
Space-charge-induced resonance crossing is one notable limitation of beam intensity in high-intensity synchrotrons. This paper proposes a modification to the Resonance Driving Terms (RDTs) to compensate for the combined effects of space charge and magnetic field imperfections under resonance crossing. The new RDTs are named modified RDTs. The effectiveness of the modified RDTs is demonstrated through simulations of half-integer and 3rd-order resonance crossings using the lattice of the High Intensity Heavy-Ion Accelerator Facility Booster Ring (HIAF-BRing). The simulations illustrate that the compensation provided by the modified RDTs significantly suppresses emittance growth and reduces distortion in the phase space.
@@ I Introduction
Space charge effect is the key limitation of beam intensity in high intensity synchrotrons. The space charge force generates an incoherent tune spread, because particles with different betatron amplitudes experience different betatron tune shifts, $\Delta\nu_{sc}(J_{x,y})$, under the space charge effect. As the beam intensity gets stronger, the tune spread becomes larger, which would eventually lead to resonance crossing. This can lead to beam instability due to the combined effect of space charge and magnetic field imperfections.
@@
Here are 4 approaches that have been proposed to mitigate this issue. Adrian Oeftiger et al. proposed a novel space charge compensation concept, pulsed electron lenses. This technique aims to mitigate the ion space charge effect through the electronic space charge effect. Simulations of the Facility for Antiproton and Ion Research (FAIR) SIS100 synchrotron demonstrated that this method could achieve up to a 30% increase in maximum beam intensity [1]. Hideaki Hotchi et al. at Facilities at Japan Proton Accelerator Research Complex (J-PARC) proposed a novel method of transverse injection painting to counteract the emittance exchange caused by Montague resonance. The beam stability was achieved by sextupole compensation under resonance crossing of 3rd-order resonance lines. However, the scheme for achieving this compensation in the rapid cycling synchrotron (J-PARC-RSC) was not demonstrated in the cited reference[2]. Meanwhile, the compensation scheme in the Main Ring (J-PARC-MR) was demonstrated but the authors considered that the optimal compensation had not been obtained through their scheme[3]. Andrea Santamaria Garcia and Foteini Asvesta et al. achieved the beam stability under resonance crossing of a half-integer resonance line by scanning currents of quadrupole correctors at Proton Synchrotron Booster[4, 5].
@@
In these work, Resonance Driving Terms (RDTs[6]) play an important role in resonance compensation. RDTs focus on identifying and quantifying the effects of specific perturbations that could induce resonant behavior of the beam. These perturbations include imperfections in the magnetic fields of the accelerator. The Hamiltonian of the system could be expressed as a series of Fourier coefficients depending on the amplitude and phase of the beam oscillations[7].
@@
The key idea of this paper is the introduction of space charge modification in RDTs. This idea was verified through simulations performed on a dedicated accelerator collective instability platform, CISP[8], using the lattice of the High Intensity heavy-ion Accelerator Facility Booster Ring (HIAF-BRing[9]). In the simulations, non-structural half-integer and 3rd-order resonance lines were crossed due to the space charge incoherent tune spread. Magnetic field errors were introduced to drive corresponding resonances. The modified RDTs were then used to calculate the strength of magnetic correctors in order to compensate for these errors. For comparison, the compensation of (unmodified) RDTs and the J-PARC-MR scheme were also demonstrated.
@@
The paper is structured as follows. In Sec. II, the space charge modification in RDTs is demonstrated. In Sec. III, schemes of modified RDTs, RDTs without modification and the J-PARC-MR scheme, are used in the simulation to compensate for corresponding magnetic field errors during resonance crossing. The half-integer and 3rd-order resonance crossing are successfully compensated by the modified RDTs. Sec. IV contains a summary of the whole paper.
II Space charge modification in RDTs
@@
where $n$ is the order of resonance, $n=j+k+l+m$, $K$ is the normal multipole coefficient, $J$ is the skew multipole coefficient and $L$ is the length of the $\omega^{th}$ multipole. The RDTs are useful tools to study the effect of magnetic nonlinear fields in the betatronic motion. In the following text, space charge would be introduced in order to study the combined effect of space charge and magnetic field imperfections.
When space charge could not be neglected, treating the space-charge potential $U_{sc}$ as a perturbation, the Hamiltonian becomes:
@@
where $\sigma_{u}$ is the envelope radii, the $K_{sc}$ is the “normalized” space-charge perveance parameter given by $K_{sc}=2\lambda r_{0}/(\beta_{rel}^{2}\gamma_{rel}^{3})$, where $\lambda$ could be obtained by the bunch intensity $N_{B}$ and the longitudinal rms bunch length $\sigma_{s}$, $\lambda=N_{B}/(\sqrt{2\pi}\sigma_{s})$, $r_{0}$ is the particle classical radius $e^{2}/(4\pi\varepsilon_{0}mc^{2})$, $\beta_{rel}$ and $\gamma_{rel}$ are Lorentz factors. For particles with small amplitude at the center of the beam, Eq. (5) could be expanded in Taylor series in order to get its linear strength:
$\frac{\partial U_{sc}}{\partial u}=-\frac{K_{sc}}{\sigma_{u}(\sigma_{x}+\sigma_{y})}u=-K_{u,sc}u+...$ (6)
@@
from which a new series of Twiss parameters influenced by space charge could be obtained, but they could only apply to particles with small amplitude. For particles with larger amplitudes, the absolute value of the defocusing quadrupole strength $-K_{u,sc}$ is smaller than that given by Eq.(6), resulting in a correspondingly smaller tune shift. This has been demonstrated by the $Poincar\acute{e}$ map in [12], the frequency map analysis in [13], and distribution density of the space-charge tune shift in [14]. In order to describe the betaron motion of all particles in a bunch, Eq. (7) could be rewritten in the form that depends on the action $J_{u}$ of particles:
@@
Twiss parameters influenced by space charge of particles with any $J_{u}$ could be obtained by analysing Eq. (5,8). However, focusing on particles undergoing resonance crossing through Eq. (8) is complex. The primary objective of this paper is to compensate for resonance crossing, necessitating a focus on specific particles whose tunes intersect resonance lines. The largest betatron tune shift for a beam with a bi-Gaussian charge distribution is given by [11]:
@@
where the $\beta_{u}$ is the original $\beta$ function without space charge. Substituting Eq. (9) back into Eq. (6), the linear strength becomes the form that depends on $\Delta\nu_{sc,u}$:
@@
which means that Twiss parameters influenced by space charge, $\beta_{u,sc}$ and phase $\phi_{u,sc}$, could be obtained by Eq. (10) through adjusting the value of $\Delta\nu_{sc,u}$. Then the Twiss of space-charge-influenced particles in the vicinity of resonance lines could be obtained. For other kinds of beam distribution, the Eq. (6,9) would change into corresponding form. But the Eq. (10) could always be established. So this approach could be applied to any beam distribution.
@@
Here is an example on how to establish the modified RDTs compensation. Assuming that the chromaticity is 0, when a low-intensity beam is injected into the ring, the tune will be centered around the bare tune, $Q_{x}=9.47,Q_{y}=9.43$ in Fig. (1). As beam intensity increases, the tune spread enlarges until it hits the resonance line $283Q_{y}=$ 28, shown as the orange tune spread in Fig. (1). The original skew sextupoles in the ring would cause a stop band at this resonance line, limiting further intensity increase. By applying Eq. (10, 11), the strength of the skew sextupole correctors can be determined to minimize the first group of modified RDTs $f_{0030}(\beta_{u,sc},\phi_{u,sc})$ of the core particles, thereby mitigating the stop band and facilitating resonance crossing at $283Q_{y}=$ 28. Then as the intensity keeps increasing, the largest tune shift would hit the resonance line $Q_{x}+2Q_{y}=28$, shown as the green tune spread in Fig. (1). By applying Eq. (10,11), the strength of the sextupole correctors can be calculated to minimize the second group of modified RDTs $f_{1020}(\beta_{u,sc},\phi_{u,sc})$ of the core particles, facilitating resonance crossing at $Q_{x}+2Q_{y}=28$. At this stage, the first group of modified RDTs $f_{0030}(\beta_{u,sc},\phi_{u,sc})$ now describes the nonlinear motion of larger-amplitude particles, rather than the core particles. This means the ignored space-charge-induced nonlinear force becomes stronger. According to the following simulation, the space-charge-induced nonlinear force would make Eq. (10) imprecise for halo particles with large amplitude at the edge of the beam[15].
@@
In a word, this approach is precise for core particles but imprecise for halo particles, as the dynamics of halo particles is complicated for the nonnegligible nonlinearity at the edge of the beam.
More examples of using Eq. (11) for compensation can be found in Sec. III.
@@ III Simulation
III.1 Half-integer resonance crossing at $192Q_{x}=$ 19
@@
For modified RDTs, firstly, the raw Twiss parameters at (9.52,9.43) could be obtained by the ’TWISS’ order of MAD-X[16], which could provide the raw $\beta_{u}$ and $\phi_{u}$. Second, since the tune spread would cross $Q_{x}=9.5$, substituting $\Delta\nu_{sc,x}=(9.5-9.52+o)$ ($o$ asymptotically approaches zero to prevent the denominator of $f_{jklm}^{(b)}$ from going to zero.) into Eq. (10):
@@
Figure 4: The horizontal emittance change under the tune spread in Fig. (2). The quadropole errors come from gaussian random values, $3\sigma(\Delta B_{1}/B_{1})=1\times 10^{-3}$, on all quadropoles. In the top subgraph, the emittance without compensation grows too much, causing that rest results almost overlap. The bottom subgraph shows more details about the comparison between these compensation schemes.
Figure 5: The normalized phase spaces under the tune spread in Fig. (2). The top left subgraph shows the phase space without errors. The top right subgraph shows that with modified RDTs. The bottom left subgraph shows that with raw RDTs. The bottom right subgraph shows that with the scheme of J-PARC-MR. In the unshown phase space without compensation, which is a mess, the amplitude of almost each particle grows rapidly.
@@
For the J-PARC-MR scheme, first match the tune into the vicinity of the half-integer resonance line (9.50+$o$,9.39) and get Twiss parameters in MAD-X, $\beta_{x}(9.501,9.395),\phi_{x}(9.501,9.395)$. Then get RDTs through this Twiss and minimise the RDTs along the ring:
@@
Finally, match the tune back to (9.52,9.43) for simulation.
@@
Fig. (4) shows the change in emittance for each compensation scheme without chromaticity. The modified RDTs compensation scheme performs the best, demonstrating small emittance growth during the first 30 turns and negligible growth in the rest of the evolution. In contrast, the simulation without compensation shows rapid emittance growth during the initial 50 turns. The emittance growth then stabilizes between 50 and 1000 turns, as the reduced tune spread, resulting from beam loss and emittance increase, no longer intersects the half-integer resonance line.
@@
Fig. (5) presents the normalized phase spaces, clearly illustrating the presence of resonance islands in the phase spaces with raw RDTs and the J-PARC-MR scheme. In contrast, the phase space with modified RDTs shows little resonance behavior. The stable resonance islands observed in the bottom two subgraphs are due to insufficient compensation. This phenomenon explains why the emittance with raw RDTs and J-PARC-MR scheme stops growing at 600-1000 turns. The phase space without compensation is not shown, as the amplitude of each particle increases rapidly, leading to chaotic particle distribution with almost no useful information.
@@
The Fig. (2-5) could be redrawn as Fig. (6) after introducing chromaticity into the simulation. As shown in the top left subgraph, chromaticity introduces an additional tune shift $\xi_{u}\delta$ to each particle, which varies with the longitudinal phase space oscillation. In the simulation, the Radio-Frequency(RF) cavity maintains a voltage of 65,000 V and a phase of 0 for the ideal particle. The synchrotron tune is approximately 0.00625. After introducing chromaticity, the emittance growth for each compensation scheme becomes larger. The modified RDTs scheme still has the best performance but its emittance also grows slowly compared to the results shown in Fig. (4). This increase may be attributed to the neglect of chromaticity effects in the modified RDTs. In the right 4 subgraphs, resonance islands are no longer present, but oscillation distortions can be observed in the bottom two subgraphs. Comparing the phase space with the modified RDTs scheme to that without errors, it can be seen that the amplitude of the particles with the modified RDTs scheme varies slightly, but the phase space remains consistent with the ideal case. This confirms the successful compensation for half-integer resonance crossing using the modified RDTs scheme.
@@
III.2 3rd-order resonance crossing
@@
Figure 8: The emittance change under the tune spread in Fig. (7). The sextupole and skew sextupole errors come from $3\sigma(\Delta B_{2}/B_{1})=1\times 10^{-3}$ of all quadropoles.
Figure 9: The normalized phase spaces under the tune spread in Fig. (7). The left 2 subgraphs show the phase space without errors. The middle 2 subgraphs show that with modified RDTs. The right 2 subgraphs show that with the scheme of J-PARC-MR.
Figure 10: The action in Fig. (9). The left 2 subgraphs show the phase space without errors. The middle 2 subgraphs show that with modified RDTs. The right 2 subgraphs show that with the scheme of J-PARC-MR.
@@
where the 5 groups of modified RDTs should be minimised by corresponding correctors. For raw RDTs, just get 5 groups of RDTs at (9.47,9.43) and minimise them. For the scheme of J-PARC-MR, match the tune to the 5 black dots in Fig. (7). Corresponding groups of RDTs should be obtained at each tune. Minimise the RDTs and match the tune back to (9.47,9.43) for simulation.
@@
Fig. (8) shows the emittance change without chromaticity of each compensation scheme. The Montague resonance[19] might have a small influence on the beam, leading to a weak action transfer from the horizontal to the vertical plane even in the absence of magnetic field errors. The compensation scheme of modified RDTs demonstrates the best performance. For modified RDTs, the $\varepsilon_{rms,y}$ grows by 2%, due to the insufficient compensation for halo particles at $283Q_{y}=$ 28. The statistical RMS emittance without compensation and with raw RDTs decreases due to beam loss.
@@
The normalized phase spaces without errors, with modified RDTs and with the scheme of J-PARC-MR are demonstrated in Fig. (9). In phase space without errors, the variable $J_{y}$ and $\nu_{y}$ of the halo particles could be observed, leading to imprecise results of Eq. (10). The envelope radii in the vertical plane exceed the designed aperture even without imperfections, resulting in 3.5% of the beam being lost during acceleration when the aperture is set to the designed (200,100) $\pi$ mm mrad. In the phase space compensated by modified RDTs, oscillation distortion could be observed in the horizontal plane but the RMS emittance remains almost the same as that without errors. A slight $283Q_{y}=$ 28 resonance behavior on the black halo particle could be observed in the vertical plane. For the phase space compensated by the J-PARC-MR scheme, an obvious $283Q_{y}=$ 28 resonance behavior on the black halo particle could be observed in the vertical plane.

### v2

@@ Abstract
Space-charge-induced resonance crossing is one notable limitation of beam intensity in high-intensity synchrotrons. Beam instability occurs under the combined effects of space-charge-induced tune spread and magnetic field imperfections prevents, which limits the intensity. This paper proposes a space-charge-Twiss modification to the Resonance Driving Terms (RDTs). The new RDTs are named modified RDTs. Modified RDTs aim to describe the nonlinear particle dynamics under resonance crossing. The feasibility of the modified RDTs is demonstrated through simulations of half-integer and 3rd-order resonance crossings, using the lattice of the High Intensity Heavy-Ion Accelerator Facility Booster Ring (HIAF-BRing). The simulations illustrate that the compensation provided by the modified RDTs significantly suppresses emittance growth and reduces distortion in the phase space.
@@ I Introduction
Space charge effect is the key limitation of beam intensity in high intensity synchrotrons. It can lead to undesired beam emittance growth, beam halo formation and particle loss. Since particles with different betatron amplitudes have different betatron tune shifts, $\Delta\nu_{sc}(J_{x,y})$, the space charge force produces a incoherent tune spread. As the beam intensity gets stronger, the tune spread becomes larger, which would eventually lead to resonance crossing. Besides, the variation of the tunes through the synchrotron motion due to the longitudinal dependency of space charge and chromatic effects leads to periodic resonance crossing. The space-charge-induced resonance crossing has been identified as the main beam degradation mechanism in Proton Synchrotron (PS) and the Super Proton Synchrotron (SPS) at CERN[1].
@@
Here are 4 approaches that have been proposed to mitigate this issue. Adrian Oeftiger et al. proposed a novel space charge compensation concept, pulsed electron lenses. This technique aims to mitigate the ion space charge effect through the electronic space charge effect. Simulations of the Facility for Antiproton and Ion Research (FAIR) SIS100 synchrotron demonstrated that this method could achieve up to a 30% increase in maximum beam intensity [2]. Hideaki Hotchi et al. at Facilities at Japan Proton Accelerator Research Complex (J-PARC) proposed a novel method of transverse injection painting to counteract the emittance exchange caused by Montague resonance. The beam stability was achieved by sextupole compensation under resonance crossing of 3rd-order resonance lines. However, the scheme for achieving this compensation in the rapid cycling synchrotron (J-PARC-RSC) was not demonstrated in the cited reference[3]. Meanwhile, the compensation scheme in the Main Ring (J-PARC-MR) was demonstrated in [4]. One scheme in this paper references the scheme of J-PARC-MR. Andrea Santamaria Garcia and Foteini Asvesta et al. achieved the beam stability under resonance crossing of a half-integer resonance line by scanning currents of quadrupole correctors at Proton Synchrotron Booster[5, 6].
@@
In these work, Resonance Driving Terms (RDTs[7]) play an important role in resonance compensation. RDTs focus on identifying and quantifying the effects of specific perturbations that could induce resonant behavior of the beam. These perturbations include imperfections in the magnetic fields of the accelerator. The Hamiltonian of the system could be expressed as a series of Fourier coefficients depending on the amplitude and phase of the beam oscillations[8]. In addition, RDTs from space charge potential were carried out in [9]. However, RDTs do not consider the influence of space charge on Twiss, which causes that RDTs do not have a good performance on compensating for magnetic field imperfections under space-charge-induced resonance crossing.
@@
The key idea of this paper is the introduction of space-charge-Twiss modification in RDTs. This idea was verified through simulations performed on a dedicated accelerator collective instability platform, CISP[10], using the lattice of the High Intensity heavy-ion Accelerator Facility Booster Ring (HIAF-BRing[11]). In the simulations, non-structural half-integer and 3rd-order resonance lines were crossed due to the space charge incoherent tune spread. Magnetic field errors were introduced to drive corresponding resonances. The modified RDTs were then used to calculate the strength of magnetic correctors in order to compensate for these errors. For comparison, 2 additional compensation schemes were also demonstrated.
@@
The paper is structured as follows. In Sec. II, the space-charge-Twiss modification in RDTs is demonstrated. In Sec. III, modified RDTs scheme and 2 additional schemes are used in the simulation to compensate for corresponding magnetic field errors during resonance crossing. The half-integer and 3rd-order resonance crossing are successfully compensated by the modified RDTs. Sec. IV contains a summary of the whole paper.
II Space-charge-Twiss modification in RDTs
@@
where $n$ is the order of resonance, $n=j+k+l+m$, $K$ is the normal multipole coefficient, $J$ is the skew multipole coefficient and $L$ is the length of the $\omega^{th}$ multipole. When space charge effect could be ignored, the RDTs are useful tools to study the effect of magnetic nonlinear fields in the betatronic motion. Fig. (1) demonstrates a simple example of 3-order resonance compensation obtained by RDTs.
Figure 1: The left subgraph shows the normalized phase space with a small sextupole error at $283Q_{x}=$ 28. The right subgraph shows the normalized phase space with compensation obtained by RDTs. This simulatio is constructed without space charge.
However, space charge is not included in RDTs. As a result, RDTs do not have a good performance on multipole-driven resonance compensation under space-charge-induced resonance crossing. In the following text, space charge would be introduced in order to study the combined effect of space charge and magnetic field imperfections.
When space charge could not be ignored, treating the space-charge potential $U_{sc}$ as a perturbation, the Hamiltonian becomes:
@@
where $\sigma_{u}$ is the envelope radii, the $K_{sc}$ is the “normalized” space-charge perveance parameter given by $K_{sc}=2\lambda r_{0}/(\beta_{rel}^{2}\gamma_{rel}^{3})$, where $\lambda$ could be obtained by the bunch intensity $N_{B}$ and the longitudinal rms bunch length $\sigma_{s}$, $\lambda=N_{B}/(\sqrt{2\pi}\sigma_{s})$, $r_{0}$ is the particle classical radius $e^{2}/(4\pi\varepsilon_{0}mc^{2})$, $\beta_{rel}$ and $\gamma_{rel}$ are Lorentz factors. For particles with small amplitude at the center of the beam, Eq. (5) could be expanded in Taylor series, the first order of which demonstrates its linear defocusing quadrupole strength:
$\frac{\partial U_{sc}}{\partial u}=-\frac{K_{sc}}{\sigma_{u}(\sigma_{x}+\sigma_{y})}u+\cdot\cdot\cdot=-K_{u,sc}u+\cdot\cdot\cdot$ (6)
@@
from which a new series of Twiss parameters influenced by space charge could be obtained, but they could only apply to particles with small amplitude. For particles with larger amplitudes, the absolute value of the defocusing quadrupole strength $-K_{u,sc}$ is smaller than Eq.(6), leading to a correspondingly smaller tune shift. This has been demonstrated by the $Poincar\acute{e}$ map in [14], the frequency map analysis in [15], and distribution density of the space-charge tune shift in [16]. In order to describe the betaron motion of all particles in a bunch, Eq. (7) could be rewritten in the form that depends on the action $J_{u}$ of particles:
@@
Twiss parameters influenced by space charge of particles with any $J_{u}$ could be obtained by Eq. (5-8). However, focusing on particles undergoing resonance crossing through Eq. (8) is complex. The primary objective of this paper is to compensate for resonance crossing, necessitating a focus on specific particles whose tunes intersect resonance lines. The largest betatron tune shift for a beam with a bi-Gaussian charge distribution is given by [13]:
@@
where the $\beta_{u}$ is the original $\beta$ function without space charge. Substituting Eq. (9) back into Eq. (6), the linear strength becomes the form that depends on the tune shift $\Delta\nu_{sc,u}$:
@@
which demonstrates that Twiss parameters influenced by space charge, $\beta_{u,sc}$ and phase $\phi_{u,sc}$, could be obtained by Eq. (10) through adjusting the value of $\Delta\nu_{sc,u}$. Then the Twiss of space-charge-influenced particles in the vicinity of resonance lines could be obtained. For other kinds of beam distribution, the Eq. (6,9) would change into corresponding form. But the Eq. (10) could always be established. So this approach could be applied to any beam distribution.
@@
Here is an example on how to establish the modified RDTs compensation. Assuming that the chromaticity is 0, when a low-intensity beam is injected into the ring, the tune will be centered around the bare tune, $Q_{x}=9.47,Q_{y}=9.43$ in Fig. (2). As beam intensity increases, the tune spread enlarges until it hits the resonance line $283Q_{y}=$ 28, shown as the orange tune spread in Fig. (2). The original skew sextupoles in the ring would cause a stop band at this resonance line, limiting further intensity increase. By applying Eq. (10, 11), the strength of the skew sextupole correctors can be determined to minimize the first group of modified RDTs $f_{0030}(\beta_{u,sc},\phi_{u,sc})$ of the core particles, thereby mitigating the stop band and facilitating resonance crossing at $283Q_{y}=$ 28. Then as the intensity keeps increasing, the largest tune shift would hit the resonance line $Q_{x}+2Q_{y}=28$, shown as the green tune spread in Fig. (2). By applying Eq. (10,11), the strength of the sextupole correctors can be calculated to minimize the second group of modified RDTs $f_{1020}(\beta_{u,sc},\phi_{u,sc})$ of the core particles, facilitating resonance crossing at $Q_{x}+2Q_{y}=28$. At this stage, the first group of modified RDTs $f_{0030}(\beta_{u,sc},\phi_{u,sc})$ now describes the nonlinear motion of larger-amplitude particles, rather than the core particles. This means the ignored space-charge-induced nonlinear force becomes stronger. According to the following simulation, the space-charge-induced nonlinear force would make Eq. (10) imprecise for halo particles[17] with large amplitude at the edge of the beam. In a word, this approach is precise for core particles but imprecise for halo particles, as the dynamics of halo particles is complicated for the nonnegligible nonlinearity at the edge of the beam. More examples of using Eq. (11) for compensation can be found in Sec. III.
@@ III Simulation
III.1 Systematic resonance analysis
III.2 Half-integer resonance crossing at $192Q_{x}=$ 19
@@
For modified RDTs, firstly, the raw Twiss parameters at (9.52,9.43) could be obtained by the ’TWISS’ order of MAD-X[19], which could provide the raw betatron function $\beta_{u}$ and phase $\phi_{u}$. Second, since the tune spread would cross $Q_{x}=9.5$, substituting $\Delta\nu_{sc,x}=(9.5-9.52+o)$ ($o$ asymptotically approaches zero to prevent the denominator of $f_{jklm}^{(b)}$ from going to zero.) into Eq. (10):
@@
For another scheme, first adjust the tune into the vicinity of the half-integer resonance line (9.50+$o$,9.39) through ’MATCH’ order of MAD-X and get Twiss parameters, $\beta_{x}(9.501,9.395),\phi_{x}(9.501,9.395)$. Then get RDTs through this Twiss and minimise the RDTs along the ring:
@@
Finally, match the tune back to (9.52,9.43) for simulation. In the later text, this scheme is called moved RDTs.
The key difference among these 3 schemes is the Twiss: modified RDTs use space-charge-influenced Twiss, raw RDTs use the Twiss at operating point (9.52,9.43), moved RDTs use the Twiss at the intersection of the tune spread and resonance lines through ‘MATCH’ order of MAD-X.
@@
Fig. (6) shows the emittance evolution for each compensation scheme without chromaticity. The modified RDTs compensation scheme performs the best, demonstrating small emittance growth during the first 30 turns and negligible growth in the rest of the evolution. In contrast, the simulation without compensation shows rapid emittance growth during the initial 50 turns. The emittance growth then stabilizes between 50 and 1000 turns, as the reduced tune spread, resulting from beam loss and emittance increase, no longer intersects the half-integer resonance line.
Figure 6: The horizontal emittance evolution under the tune spread in Fig. (4). The quadropole errors come from gaussian random values, $3\sigma(\Delta B_{1}/B_{1})=1\times 10^{-3}$, on all quadropoles. In the top subgraph, the emittance without compensation grows too much, causing that rest results almost overlap. The bottom subgraph shows more details about the comparison between these compensation schemes.
Figure 7: The normalized phase spaces under the tune spread in Fig. (4). The top left subgraph shows the phase space without errors. The top right subgraph shows that with modified RDTs. The bottom left subgraph shows that with raw RDTs. The bottom right subgraph shows that with moved RDTs. In the unshown phase space without compensation, which is a mess, the amplitude of almost each particle grows rapidly.
@@
Fig. (7) presents the normalized phase spaces, clearly illustrating the presence of resonance islands in the phase spaces with raw RDTs and moved RDTs. In contrast, the phase space with modified RDTs shows little resonance behavior. The stable resonance islands observed in the bottom two subgraphs are due to insufficient compensation. This phenomenon explains why the emittance with raw RDTs and moved RDTs stops growing at 600-1000 turns. The phase space without compensation is not shown, as the amplitude of each particle increases rapidly, leading to chaotic particle distribution with almost no useful information.
@@
The Fig. (4-7) could be redrawn as Fig. (8) after introducing chromaticity into the simulation. As shown in the top left subgraph, chromaticity introduces an additional tune shift $\xi_{u}\delta$ to each particle, which varies with the longitudinal phase space oscillation. In the simulation, the Radio-Frequency(RF) cavity maintains a voltage of 65,000 V and a phase of 0 for the ideal particle. The synchrotron tune is approximately 0.00625. After introducing chromaticity, the emittance growth for each compensation scheme becomes larger. The modified RDTs scheme still has the best performance but its emittance also grows slowly compared to the results shown in Fig. (6). This increase may be attributed to the neglect of chromaticity effects in the modified RDTs. In the right 4 subgraphs, resonance islands are no longer present, but oscillation distortions can be observed in the bottom two subgraphs. Comparing the phase space with the modified RDTs scheme to that without errors, it can be seen that the amplitude of the particles with the modified RDTs scheme varies slightly, but the phase space remains consistent with the ideal case. This confirms the successful compensation for half-integer resonance crossing using the modified RDTs scheme.
@@
III.3 3rd-order resonance crossing
@@
Figure 10: The emittance evolution under the tune spread in Fig. (9). The sextupole and skew sextupole errors come from $3\sigma(\Delta B_{2}/B_{1})=1\times 10^{-3}$ of all quadropoles.
Figure 11: The normalized phase spaces under the tune spread in Fig. (9). The left 2 subgraphs show the phase space without errors. The middle 2 subgraphs show that with modified RDTs. The right 2 subgraphs show that with moved RDTs.
Figure 12: The action in Fig. (11). The left 2 subgraphs show the phase space without errors. The middle 2 subgraphs show that with modified RDTs. The right 2 subgraphs show that with moved RDTs.
@@
where the 5 groups of modified RDTs should be minimised by corresponding correctors. For raw RDTs, just get 5 groups of RDTs at (9.47,9.43) and minimise them. For moved RDTs, match the tune to the 5 black dots in Fig. (9). Corresponding groups of RDTs should be obtained at each tune. Minimise the RDTs and match the tune back to (9.47,9.43) for simulation.
@@
Fig. (10) shows the emittance evolution without chromaticity of each compensation scheme. The Montague resonance[22] has a small influence on the beam, leading to a weak action transfer from the horizontal to the vertical plane even in the absence of magnetic field errors. The compensation scheme of modified RDTs demonstrates the best performance. For modified RDTs, the $\varepsilon_{rms,x}$ keeps the same as that without errors, and the $\varepsilon_{rms,y}$ grows by 2%, due to the insufficient compensation for halo particles at $283Q_{y}=$ 28. It can be observed that the statistical RMS emittance without compensation and with raw RDTs increases in the initial 20 turns, and decreases in the rest of the evolution. In fact the emittance of the beam keeps increasing, but particles at edge of the beam reach the simulation aperture, leading to the decrease of statistical RMS emittance.
@@
The normalized phase spaces without errors, with modified RDTs and with moved RDTs are demonstrated in Fig. (11). In phase space without errors, the variable $J_{y}$ and $\nu_{y}$ of the halo particles could be observed, leading to imprecise results of Eq. (10). The vertical oscillation of large amplitude particle is magnified due to space charge. As a result, the envelope radii in the vertical plane exceed the designed aperture even without imperfections, leading to 3.5% of the beam being lost during acceleration when the aperture is set to the designed (200,100) $\pi$ mm mrad. In the phase space with modified RDTs, oscillation distortion could be observed in the horizontal plane but the RMS emittance remains almost the same as that without errors. A slight $283Q_{y}=$ 28 resonance behavior on the black halo particle could be observed in the vertical plane, demonstrating the imprecise compensation for halo particles with modified RDTs. In the phase space with moved RDTs, the amplitude growth of the grey halo particle, the serious oscillation distortion of black halo particle in the horizontal plane and an obvious $283Q_{y}=$ 28 resonance behavior of the black halo particle in the vertical plane could be observed. Note that the black halo particle in the horizontal plane and the black halo particle in the vertical plane are not the same one particle.